### v1

# RestoreVAR: Visual Autoregressive Generation for All-in-One Image Restoration

Abstract
The use of latent diffusion models (LDMs) such as Stable Diffusion has significantly improved the perceptual quality of All-in-One image Restoration (AiOR) methods, while also enhancing their generalization capabilities. However, these LDM-based frameworks suffer from slow inference due to their iterative denoising process, rendering them impractical for time-sensitive applications. To address this, we propose RestoreVAR, a novel generative approach for AiOR that significantly outperforms LDM-based models in restoration performance while achieving over $\mathbf{10\times}$ faster inference. RestoreVAR leverages visual autoregressive modeling (VAR), a recently introduced approach which performs scale-space autoregression for image generation. VAR achieves comparable performance to that of state-of-the-art diffusion transformers with drastically reduced computational costs. To optimally exploit these advantages of VAR for AiOR, we propose architectural modifications and improvements, including intricately designed cross-attention mechanisms and a latent-space refinement module, tailored for the AiOR task. Extensive experiments show that RestoreVAR achieves state-of-the-art performance among generative AiOR methods, while also exhibiting strong generalization capabilities. Project page: https://sudraj2002.github.io/restorevarpage/
1 Introduction
Image restoration is a complex inverse problem that aims to recover clean images from various forms of degradations, such as haze, rain, snow, blur, and low-light conditions. Recently, the paradigm of All-in-One image Restoration (AiOR) has emerged, where a single deep network is trained to handle multiple degradation types. Existing AiOR methods can be broadly categorized into non-generative and generative approaches. Non-generative models such as AirNet airnet , PromptIR promptir , InstructIR instructir , AWRaCLe awracle , and AdaIR adair , deterministically map degraded images to their clean counterparts. While these methods offer fast inference and reliable pixel-level restoration performance, they often fail to generalize to diverse degradations encountered in real-world scenarios. To overcome this challenge, recent works have adopted generative models that aim to capture the distribution of clean images and produce more perceptually realistic outputs. Early works gan1rain ; gan2rain ; gan3blur based on GANs gan attempted this through adversarial learning, but suffered from mode collapse and unstable training. To improve fidelity and training stability, methods such as DiffUIR diffuir and DA-CLIP daclip employed pixel-space diffusion models ddpm . However, their high computational cost makes large-scale pretraining infeasible, limiting their ability to learn strong generative priors. In contrast, recent methods such as Diff-Plugindiffplugin , AutoDIRautodir , and PixWizard pixwizard leverage latent diffusion models (LDMs), such as Stable Diffusion stablediff . By operating in a lower-dimensional latent space, LDMs significantly reduce computational costs, enabling large-scale pretraining which equips them with strong generative priors of natural images. These priors allow LDM-based AiOR methods to deliver perceptually realistic restoration and improved generalization to real-world degradations.
Despite their advantages, LDM-based AiOR methods have some shortcomings. (1) LDMs require multiple denoising steps during inference, resulting in significantly longer runtimes compared to non-generative models. Their slow inference speeds pose challenges for applications that demand real-time processing, such as video surveillance or autonomous navigation. (2) LDMs rely on variational autoencoders (VAEs) vae which are primarily trained for generative diversity, rather than accurate pixel-level reconstruction. Consequently, the restored images obtained from LDM-based AiOR methods exhibit loss of fine structural details, hindering their performance.
An alternative class of generative models are autoregressive models, which have driven rapid advances in natural language processing through large language models (LLMs) such as GPT-3 radford2019language and LLaMA touvron2023llama . These models generate outputs by predicting the next token in a sequence, conditioned on previously generated tokens. Recently, Visual AutoRegressive (VAR) Modeling var introduced a scale-space autoregressive framework for image generation, performing next-scale prediction in the latent space of a multi-scale vector-quantized VAE (VQVAE). VAR achieves performance comparable to those of state-of-the-art diffusion models such as DiT-XL/2 dit , while operating 45×45$\times$45 × faster. Despite its success in generative tasks, the application of VAR to low-level vision tasks such as image restoration remains largely unexplored. To the best of our knowledge, only two prior works—VarSR varsr and Varformer varformer —have attempted to use VAR for image restoration. VarSR focused exclusively on the super-resolution task, while Varformer utilized intermediate VAR features to guide a separate non-generative network for AiOR. In contrast, our approach is generative and fully exploits the strong priors of the pretrained VAR model by training it directly for the AiOR task.
(a) Overview
(b) PSNR (dB)$\uparrow$
(c) Model complexity (bubble size $\propto$ parameter count)
Figure 1: RestoreVAR, our proposed VAR-based var scale-space generative AiOR model (a), significantly outperforms LDM-based methods as shown in (b). RestoreVAR also offers drastic reductions in computational complexity as shown in (c).
To this end, we introduce RestoreVAR, a novel generative approach for AiOR that addresses some of the key limitations of LDM-based approaches. Firstly, RestoreVAR adopts the autoregressive structure of VAR, achieving state-of-the-art generative AiOR performance with over $\mathbf{10\times}$ faster inference than LDM-based methods (see Fig. 1). Secondly, RestoreVAR employs cross-attention mechanisms conditioned on the degraded image latents, enabling the model to maintain spatial consistency and minimize hallucinations. Thirdly, to mitigate the loss of fine details by the vector quantization and VAE decoding processes, we propose a lightweight (only $\sim\mathbf{3}$overhead) non-generative latent refinement transformer which predicts de-quantized latents from the outputs of VAR. Additionally, we fine-tune the VAE decoder to operate on these continuous latents, further enhancing reconstruction quality. Finally, through extensive experiments, we demonstrate that RestoreVAR achieves state-of-the-art performance among generative restoration models, while also exhibiting strong generalization to real-world degradations. To summarize, our key contributions are:
  1. 1.
    We propose RestoreVAR, a novel generative AiOR framework that achieves superior performance and a $\mathbf{10\times}$ faster inference than LDM-based methods. To the best of our knowledge, RestoreVAR is the first method to directly utilize the VAR model for the AiOR task.
  2. 2.
    To achieve semantically coherent restoration, we introduce degraded image conditioning through cross-attention at each block of the VAR transformer.
  3. 3.
    To mitigate the loss of fine details in the vector quantization and VAE decoding processes, we introduce a non-generative latent refiner transformer which converts discretized latents into continuous ones, and fine-tune the VAE decoder to operate on continuous latents.
  4. 4.
    Extensive experiments show that RestoreVAR attains state-of-the-art performance among generative AiOR approaches, with perceptually preferable results and strong generalization.
2 Related Works
We now discuss prior works on image restoration and autoregressive models in computer vision.
2.1 Image restoration
Early restoration models primarily addressed specific degradations  early1 ; haze1 ; early3 ; rain1 ; rainnew1 ; rainnew2 ; snow1 ; snow2 ; gopro ; hide . Later methods such as Restormer restormer , MPRNet mprnet and SwinIR swinir introduced baseline architectures for any single restoration task. However, they are restricted to handle one degradation at a time, making them ineffective for multiple degradations. All-in-One image Restoration (AiOR) methods aim to tackle multiple corruptions with a single model. Early approaches include non-generative models such as All-in-one nas , which used neural architecture search, and Transweather transw for adverse weather restoration. PromptIR promptir used learnable prompts while AWRaCLe awracle utilized visual in-context learning to extract degradation characteristics. Other approaches such as InstructIR instructir adopted textual guidance, and DCPT dcpt proposed a novel pre-training strategy for AiOR. Recent AiOR methods have adopted diffusion models. Pixel-space diffusion models (PSDMs) such as DA-CLIP daclip and DiffUIR diffuir demonstrated improved AiOR performance. However, PSDMs lack the robust generative priors of latent diffusion models (LDMs) such as Stable Diffusion stablediff . Thus, recent methods have utilized the strong priors of LDMs for AiOR. Diff-Plugin diffplugin adopts task plugins to guide an LDM for AiOR. AutoDIR autodir developed an automatic approach for degradation detection and restoration using an LDM. PixWizard pixwizard is a multi-task SD-XL sdxl based model capable of performing AiOR among other tasks. However, LDM-based approaches are slow at inference time—a limitation we aim to overcome using visual autoregressive modeling (VAR) var .
2.2 Autoregressive models in vision
Autoregressive (AR) models generate sequences by predicting the next element conditioned on preceding ones, modeling the joint probability of the sequence as a product of conditional probabilities. This paradigm has revolutionized natural language processing (NLP) and forms the foundation of large language models (LLMs) such as GPT-3 radford2019language and LLaMA touvron2023llama . Recent works van2016pixel ; var have extended AR models to vision and can be categorized as pixel-space AR van2016pixel ; van2016conditional ; reed2017parallel ; chen2019pixelsnail , token-based AR oord2018neural ; yu2023video ; nash2021generating ; ramesh2021zeroshot and scale-space AR var ; ren2024mvar ; ren2024flowar ; guo2025fastvar . Pixel-space AR predicts raw pixels one by one in raster order, as in PixelRNN van2016pixel and PixelCNN++ salimans2017pixelcnn++ , but is prohibitively slow at high resolutions. Token-based AR compresses images into discrete latent codes via vector quantization (e.g., VQ-VAE oord2018neural , VQ-VAE-2 razavi2019generating , VQGAN esser2021taming ) and then models code sequences with transformers (e.g., Image Transformer parmar2018image , ImageGPT chen2020image ). This trades-off codebook size and transformer capacity against tractability for high-resolution generation. Scale-space AR, as introduced in VAR var , generates latents hierarchically from coarse to fine scales and matches the quality of Diffusion Transformers dit at a fraction of the inference costvar . HART hart scales VAR to $1024\times 1024$ synthesis by using a lightweight MLP-based diffusion refiner to convert discrete VAR latents into continuous representations. Despite VAR’s success in generative tasks, it remains underexplored for image restoration. Only two prior works-VarSR varsr and Varformer varformer -have exploited VAR’s generative priors for image restoration. VarSR addressed super-resolution, while Varformer used VAR’s features to guide a non-generative AiOR model. In contrast, RestoreVAR is a generative model which directly trains VAR for multiple image restoration tasks.
Figure 2: Illustration of RestoreVAR for training and inference. (a) Shows the training procedure for each component of RestoreVAR, and (b) shows the overall pipeline during inference.
3 Proposed Method
In this section, we begin by explaining the working principles behind VAR for image generation. We then detail RestoreVAR, our proposed VAR-based approach for AiOR.
3.1 Preliminaries: Visual Autoregressive Modelling
Visual Autoregressive Modelling, or VAR, is a novel autoregressive class-conditioned image generation method which uses a GPT-2 gpt2 style decoder-only transformer architecture for next-scale prediction. The VAR transformer operates in the latent space of a multi-scale VQVAE which uses $K$ scales. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the VQVAE encoder outputs a latent representation $f_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. Hereafter, we will refer to $f_{\text{cont}}$ as the continuous latent, and the latent obtained after quantization as discrete latent. Instead of directly quantizing $f_{\text{cont}}$, a multi-scale residual quantization using a shared codebook across $K$ spatial scales is performed. First, the residual and accumulated quantized (or discrete) reconstruction of $f_{\text{cont}}$ are initialized as $f_{\text{res}}^{(0)}:=f_{\text{cont}}$ and $f_{\text{quant}}^{(0)}:=0$, respectively. At each scale $k=1,\ldots,K$, an index map $r_{k}\in\mathbb{Z}^{H_{k}\times W_{k}}$ is obtained by quantizing the downsampled residual feature:
$r_{k}:=\text{quantize}\left(\text{downsample}\left(f_{\text{res}}^{(k-1)}\right)\right).$
The indices $r_{k}$ are then decoded using the codebook embeddings $e(\cdot)$, upsampled to match the full resolution, and refined using a convolutional module $\phi_{k}(\cdot)$ to obtain
$h_{k}:=\phi_{k}\left(\text{upsample}\left(e(r_{k})\right)\right),\in\mathbb{R}^{H_{K}\times W_{K}\times C}.$
This is done to approximate the information captured at the current scale which is used to update the residual continuous features to be modelled by subsequent scales as
$f_{\text{quant}}^{(k)}:=f_{\text{quant}}^{(k-1)}+h_{k},\quad f_{\text{res}}^{(k)}:=f_{\text{cont}}-f_{\text{quant}}^{(k)}.$
This process is repeated for all scales and yields a set of index maps $\{r_{1},r_{2},\ldots,r_{K}\}$, each consisting of the code-book indices of residual information at an increasingly finer scale.
For training, VAR uses teacher-forcing, where the ground-truth index maps $\{r_{1},r_{2},\ldots,r_{K}\}$ are used to autoregressively predict the next scale. For each scale $k$, the accumulated reconstruction $f_{\text{quant}}^{(k-1)}=\sum_{i=1}^{k-1}\phi_{i}\left(\text{upsample}(e(r_{i}))\right)$ is interpolated to the resolution of scale $k$ to obtain $\hat{f}_{\text{quant}}^{(k)}$, which is then flattened into tokens, and concatenated with the remaining tokens to form the input sequence. A start-of-sequence (SOS) token, derived from the class label embedding, is then prepended to this input sequence. A block-wise causal attention mask is used to ensure that predictions for scale $k$ attend only to the previous scales. VAR is trained to minimize the cross-entropy loss between predicted logits and the ground-truth index maps, modeling the likelihood
$p(r_{1},r_{2},\ldots,r_{K})=\prod_{k=1}^{K}p(r_{k}\mid r_{1},r_{2},\ldots,r_{k-1}).$
During inference, the SOS token is created from the target class label. VAR then autoregressively predicts each index map $r_{k}$, one scale at a time. After predicting $r_{k}$, its embedding is upsampled, refined and accumulated to form the input for the next scale, mimicking the same procedure used during training. The VAR model uses only $K=10$ latent scales with key-value (KV) caching, enabling significantly faster inference compared to latent diffusion models.
3.2 RestoreVAR
We now describe RestoreVAR, our proposed approach that adapts the generative capabilities of VAR for AiOR, while leveraging its substantial inference speed advantage over LDMs. Given a degraded image $I_{\text{deg}}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict a clean output $I_{\text{clean}}$, close to the ground-truth $I_{\text{gt}}$. Adapting VAR to AiOR is non-trivial due to the need for high-quality pixel-level reconstruction, which is compromised by two factors: (1) VAR’s strong generative priors can cause hallucinations in the restored images without proper conditioning. (2) Vector quantization and VAE decoding introduce artifacts that hinder pixel-level restoration. RestoreVAR addresses these challenges through architectural enhancements, including cross-attention to incorporate semantic guidance from the degraded image, and a novel non-generative transformer that refines discrete latents into their continuous form to preserve fine details in the restored image. We describe these components below.
3.2.1 Autoregressive Transformer Architecture
For training, the multi-scale teacher-forcing input is constructed from the ground-truth image $I_{\text{gt}}$ (see Sec. 3.1). The start-of-sequence (SOS) token is computed from a fixed label index and augmented with a global context vector derived from the degraded image (see supplementary for details). These features are flattened and concatenated into a token sequence $\hat{f}_{\text{quant}}\in\mathbb{R}^{L\times C}$, where $L$ is the total number of tokens across all scales (see Fig. 2(a)(i)). The VAR transformer is then trained to autoregressively predict the next-scale indices $\{r^{k}_{\text{gt}}\}_{k=1}^{K}\in\mathbb{R}^{L}$ of the clean image.
To enable semantically consistent restoration, we inject information from the degraded image through cross-attention at each transformer block. At block $i$, the queries are given by the output of the feed-forward network ($x_{\text{block}_{i}}\in\mathbb{R}^{L\times D}$, where $D$ is the embedding dimension), while keys and values are derived from the continuous latent of the degraded image, $f_{\text{cont}}^{\text{deg}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. This latent is reshaped into a sequence of conditioning tokens and is appropriately projected to the embedding dimension of the transformer. As shown in Sec. 4.4, conditioning on continuous latents significantly outperforms conditioning on discrete ones. To summarize, cross-attention ($\text{CA}(\cdot,\cdot)$) is applied as
$x_{\text{block}_{\text{CA}}}=x_{\text{block}_{i}}+g_{i}\times\text{CA}(x_{\text{block}_{i}},f_{\text{cont}}^{\text{deg}}).$
We initialize $g_{i}=0$ to retain VAR’s pretrained behavior and gradually introduce conditioning. Furthermore, we replace absolute positional embeddings in VAR with 2D Rotary Positional Embeddings (RoPE) for scaling resolution from $256\times 256$ to $512\times 512$, as RoPE is well-suited for handling varying sequence lengths rope . We also remove AdaLN layers, reducing $\sim 100$M parameters with negligible impact on performance. Inference closely follows that of VAR (see Sec. 3.1), except that each scale prediction is now guided by the degraded latent. The output is a sequence of predicted indices $\{\hat{r}^{k}_{\text{pred}}\}_{k=1}^{K}$, which is then used to construct the discrete restored latent $f^{\text{pred}}_{\text{quant}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. The above steps are shown in Fig. 2(a)(ii). More architectural details are given in the supplementary.
3.2.2 Detail-Preserving Restoration
The discrete latent ($f^{\text{pred}}_{\text{quant}}$) predicted by the RestoreVAR transformer is decoded by the VQVAE to produce the restored image. However, vector-quantization and VAE decoding cause a noticeable loss of fine details in the pixel-space, leading to distorted reconstructions (see Fig. 4(b)). This presents a major challenge for using VAR in AiOR, as the scene semantics may not be accurately preserved. To address this, we introduce VAE decoder fine-tuning on continuous latents, and a lightweight latent refinement transformer (LRT) that converts discrete latents to continuous latents for decoding.
VAE Decoder Fine-Tuning. HART hart addressed VAE-induced distortions by fine-tuning the VAE decoder on both discrete and continuous latents. While effective for generative tasks, the VAE decoder of HART produces overly textured outputs, compromising accurate reconstruction (see Fig. 4(c)). Instead, we fine-tune the decoder only on continuous latents, bypassing the quantizer. The encoder and quantizer are kept frozen, and the decoder is trained on $(f^{\text{gt}}_{\text{cont}},I_{\text{gt}})$ pairs. To avoid overly smooth outputs, we use a PatchGAN patchgan discriminator (see Sec. 4.4) and optimize the decoder using pixel-wise, perceptual, and adversarial losses as
$\mathcal{L}_{\text{dec}}=\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{percep}}+\lambda_{4}\mathcal{L}_{\text{adv}},$
(a) Input
(b) VAR
(c) HART
(d) Ours
Figure 3: Qualitative comparisons of the input reconstructed using VAR var , HART hart and Our VAE decoder. Our result has minimal distortions.
Input
Discrete
Refiner
Continuous
Figure 4: Illustration of images decoded from discrete and continuous latents, along with the refiner’s predicted residuals.
where $\mathcal{L}_{\text{L1}}$ is the L1 loss, $\mathcal{L}_{\text{SSIM}}$ is the SSIM loss, $\mathcal{L}_{\text{percep}}$ is the perceptual loss, $\mathcal{L}_{\text{adv}}$ is the adversarial loss and $\lambda_{i}$ are their respective weights (see Fig. 2(a)(iv)). Our fine-tuning approach yields a decoder that is well-aligned with the objectives of AiOR, achieving mean (over $1000$ samples) reconstruction PSNR/SSIM scores of $28.14$dB/$0.842$, outperforming both the VAR VQVAE ($22.59$dB/$0.679$) and HART decoders ($26.48$dB/$0.804$). Fig. 4(d) shows that the reconstruction using our decoder is significantly better than VAR and HART (see zoomed patch).
Refining Discrete Latents. Since the VAE decoder is fine-tuned for continuous latents, the predicted discrete latent, $f^{\text{pred}}_{\text{quant}}$, must be converted into a continuous form for decoding. While HART uses a $37$M parameter diffusion-based MLP for this, it incurs a $\sim 20\%$ inference overhead due to iterative denoising. Instead, we propose a lightweight, non-generative latent refinement transformer (LRT) that predicts a residual, which when added to $f^{\text{pred}}_{\text{quant}}$, produces a continuous latent, $\hat{f}_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$ as
$\hat{f}_{\text{cont}}=f^{\text{pred}}_{\text{quant}}+\text{LRM}(f^{\text{pred}}_{\text{quant}},z),$
where $z\in\mathbb{R}^{L\times D}$ is the output from the final RestoreVAR transformer block. $z$ is passed through cross-attention and provides pseudo-continuous guidance to the LRT which is critical for performance (see Sec. 4.4). The LRT is trained using $\mathcal{L}_{1}$ loss between the predicted and ground-truth continuous latents ($f_{\text{cont}}^{\text{gt}}$) as $\mathcal{L}_{\text{LRT}}=\mathcal{L}_{1}(\hat{f}_{\text{cont}},f_{\text{cont}}^{\text{gt}})$. Our LRT introduces only $3\%$ additional overhead and significantly outperforms HART’s refiner in PSNR and SSIM scores (see Sec. 4.4). The training procedure of the LRT is shown in Fig. 2(a)(iii) and Fig. 4 provides a visual example of its predictions.
Thus, RestoreVAR combines the VAR transformer, LRT, and fine-tuned decoder to deliver fast, perceptually realistic, and structurally faithful results. Fig. 2(b) depicts inference of RestoreVAR.
4 Experiments
In this section, we provide implementation details, comparisons with existing All-in-One image Restoration (AiOR) approaches, and present ablations on key components of our framework.
4.1 Implementation Details
Each component of RestoreVAR was trained independently to disentangle learning objectives. We used the VAR model of depth $16$ as the RestoreVAR transformer backbone and trained it with the AdamW optimizer adamw , a learning rate (LR) of $10^{-4}$, batch size of $48$, for $100$ epochs. The latent refiner was trained for $100$ epochs with the AdamW optimizer, LR$=10^{-4}$ and a batch size of $96$. The VAE decoder was fine-tuned using a weighted combination of losses (see Sec. 3.2.2) with empirically chosen weights: $\lambda_{1}=2.0$, $\lambda_{2}=0.4$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.01$. Fine-tuning was performed for $5$ epochs with a learning rate of $3\times 10^{-4}$ and a batch size of $12$, using AdamW. All training was conducted on $8$ RTX A6000 GPUs, while inference was performed on a single RTX 4090 GPU.
4.2 Datasets
We trained RestoreVAR for five restoration tasks: dehazing, desnowing, deraining, low-light enhancement and deblurring. For dehazing, we used the RESIDE reside dataset comprising $72135$ training and $500$ test images. The Snow100k dataset snow100k was used for desnowing, with 50,000 training images and 16,801 test images (heavy subset). For deraining, we used Rain13K mprnet consisting of 13,711 training images and 4,298 test images. The LOLv1 lolv1 dataset was used for low-light enhancement, consisting of $485$ training images and $15$ test images. For deblurring, we used the GoPro gopro dataset comprising 2,103 training images and 1,111 test images. We additionally test the generalization capability of our model on real-world datasets, namely, LHP lhprain ($1000$ images), REVIDE revide ($284$ images), SICE sice ($565$ images), TOLED toledpoled ($30$ images) and POLED toledpoled ($30$ images).
4.3 Comparisons
We compare RestoreVAR with recent state-of-the-art generative and non-generative methods for AiOR. For non-generative approaches, we include PromptIR promptir , InstructIR instructir , AWRaCLe awracle and DCPT dcpt . Among generative methods, we compare with the latent diffusion model (LDM) based approaches Diff-Plugin diffplugin , AutoDIR autodir and PixWizard pixwizard . To ensure a fair comparison, we retrained PromptIR and AWRaCLe, as their official checkpoints were not trained for all our AiOR tasks. All other methods were evaluated using their publicly released checkpoints. We were unable to re-train InstructIR and DCPT as they have not released their training codes. For AutoDIR, we report results without the structure correction module, as this module functions as an independent, non-generative restoration network (more details in supplementary). The results reported for PixWizard were obtained using its publicly released checkpoint. We do not compare with task-specific restoration models, as RestoreVAR is proposed for the AiOR setting.
Table 1 presents quantitative results (PSNR, SSIM and LPIPS scores) on the RESIDE, Snow100k, Rain13K, LOLv1 and GoPro datasets. RestoreVAR significantly outperforms LDM-based AiOR methods across all metrics. These improvements are achieved at a fraction of the computational cost of LDM-based approaches (see Table 5), highlighting the efficacy of our framework. Qualitative comparisons with LDM-based methods in Fig. 5 further illustrate that RestoreVAR produces restored images of high quality while better preserving fine details. Visual results for the Snow100k and LOLv1 datasets are provided in the supplementary. While non-generative methods achieve better scores, it is important to recognize that the performance of RestoreVAR is inherently influenced by the quality of the VAE decoder; a limitation shared by all latent generative approaches. Despite this constraint, RestoreVAR narrows the gap with non-generative methods while maintaining the benefits of a generative framework, i.e., perceptually realistic results and strong generalization capabilities. To demonstrate these strengths, we evaluate generalization using no-reference image quality metrics (following prior works diffplugin ; autodir ; awracle ), and assess perceptual realism through a user study.
Table 1: Quantitative comparisons of RestoreVAR with the state-of-the-art LDM-based generative AiOR approaches, and non-generative methods. RestoreVAR significantly outperforms generative methods on PSNR, SSIM and LPIPS scores. The best generative approach is indicated in bold.
Method Venue RESIDE reside Snow100k snow100k Rain13K mprnet LOLv1 lolv1 GoPro gopro
[…]
Non-generative methods
PromptIR promptir NeurIPS’23 32.02 0.952 0.013 31.98 0.924 0.115 29.56 0.888 0.087 22.89 0.847 0.296 27.21 0.817 0.250
InstructIR instructir ECCV’24 26.90 0.952 0.017 29.56 0.885 0.088 22.81 0.836 0.132 28.26 0.870 0.146
AWRaCLe awracle AAAI’25 30.81 0.979 0.013 30.56 0.904 0.088 31.26 0.908 0.068 21.04 0.818 0.146 26.78 0.820 0.248
DCPT dcpt ICLR’25 29.10 0.968 0.017 24.11 0.766 0.203 23.67 0.863 0.106 27.92 0.877 0.169
Generative methods
Diff-Plugin diffplugin CVPR’24 23.23 0.765 0.091 21.02 0.611 0.196 21.71 0.617 0.169 19.38 0.713 0.195 21.76 0.633 0.217
AutoDIR autodir ECCV’24 24.48 0.780 0.081 19.00 0.515 0.347 23.02 0.642 0.162 19.43 0.766 0.135 23.55 0.700 0.168
PixWizard pixwizard ICLR’25 21.28 0.738 0.142 21.24 0.594 0.206 21.38 0.596 0.180 15.84 0.629 0.305 20.49 0.602 0.223
RestoreVAR (Ours) 24.67 0.821 0.074 24.05 0.713 0.156 23.97 0.700 0.153 21.72 0.782 0.126 23.96 0.737 0.167
Input Diff-Plugin diffplugin AutoDir autodir PixWizard pixwizard RestoreVAR GT
RESIDE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rain13K
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GoPro
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparisons of RestoreVAR with LDM-based generative AiOR approaches. RestoreVAR achieves consistent restoration with enhanced preservation of fine-details.
Table 2: Quantitative comparisons of generalization of RestoreVAR against state-of-the-art non-generative approaches on real-world images. The best result is indicated in bold.
Method Venue LHP lhprain REVIDE revide SICE sice TOLED toledpoled POLED toledpoled Average
[…]
PromptIR promptir NeurIPS’23 55.813 0.334 60.277 0.390 57.651 0.512 43.320 0.249 34.604 0.259 50.333 0.349
InstructIR instructir ECCV’24 58.288 0.335 63.103 0.352 66.089 0.493 44.985 0.298 23.317 0.241 51.156 0.344
AWRaCLe awracle AAAI’25 56.767 0.361 58.295 0.351 62.067 0.512 45.932 0.310 39.493 0.264 52.511 0.360
DCPT dcpt ICLR’25 58.075 0.341 60.109 0.389 65.293 0.521 44.298 0.281 38.241 0.323 53.203 0.371
RestoreVAR 57.662 0.414 63.607 0.429 66.147 0.537 52.374 0.338 48.118 0.276 57.582 0.399
Input PromptIR promptir InstructIR instructir AWRaCLe awracle DCPT dcpt RestoreVAR GT
LHP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
REVIDE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SICE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
TOLED
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
POLED
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative comparisons of RestoreVAR with non-generative methods on real degradations. RestoreVAR consistently achieves better results, demonstrating its generalization capability.
For testing generalization, we report MUSIQ musiq and CLIPIQA clipiqa scores in Table 2 on the real-world datasets discussed in Sec. 4.2. RestoreVAR achieves higher scores than non-generative models (on average), indicating better robustness under real degradations. Qualitative results for this experiment are shown in Fig. 6, where RestoreVAR consistently outperforms non-generative approaches.
Table 3: Mean scores from user study.
[…]
PromptIR promptir 2.05
InstructIR instructir 2.71
AWRaCLe awracle 2.11
DCPT dcpt 2.23
AutoDIR autodir 3.80
RestoreVAR 4.42
To further evaluate perceptual quality, we conducted a user study in which participants rated outputs from non-generative models, AutoDIR (LDM-based) and RestoreVAR, for $40$ real-world scenes. We received $32$ responses with each participant scoring outputs based on scene consistency, restoration quality, and overall appeal on a 5-point scale. Table 3 shows that RestoreVAR received the highest average ratings (across all three criteria), highlighting its ability to produce images that align closely with human preferences. More visual results are given in the supplementary.
Computational Complexity. RestoreVAR achieves substantial performance improvements over LDM-based AiOR approaches at a fraction of their computational cost. To show this, we compared RestoreVAR with Diff-Plugin, AutoDIR, and PixWizard in terms of inference steps, runtime, TeraFLOPs, and total parameter count. As shown in Table 5, RestoreVAR achieves a $\mathbf{10\times}$ speed-up over Diff-Plugin and a $\sim\mathbf{16\times}$ reduction in TFLOPs. Compared to AutoDIR and PixWizard, RestoreVAR is over $\mathbf{40\times}$ faster in inference.
These results highlight the effectiveness of RestoreVAR’s generative capabilities while being significantly faster than LDM-based methods.
Limitations and Scope for Future Work. Despite the strengths of RestoreVAR, there remains scope for improvement. First, its performance is inherently constrained by the latent refiner transformer (LRT) and the VAE decoder. While the LRT significantly improves results over using no refiner, it does not reach the upper bound set by directly decoding from ground-truth continuous latents. Exploring improved VQVAE and refiner architectures could help address this. Another promising direction is to employ our non-generative LRT in fully generative VAR models, given its strong performance for AiOR. Finally, future work can investigate how the performance of RestoreVAR scales with larger VAR backbones.
Table 4: Comparison of the computational complexity of RestoreVAR with LDM-based AiOR approaches.
Method Steps Time (s) TFLOPs Params (M)
Diff-Plugin 20 2.04 16.08 859.50
AutoDIR 100 8.477 67.80 859.50
PixWizard 60 8.247 19.27 2011.40
RestoreVAR 10 0.201 1.05 296.95
Table 5: Ablations on the types of latent refiners. Our proposed latent refiner transformer (LRT) performs best, with minimal overhead.
Refiner Type Time (s) Params (M) PSNR / SSIM
No Refiner 21.71 / 0.690
HART Refiner 0.0455 36.06 23.48 / 0.777
LRT w/o Last-Block 0.0036 14.61 21.23 / 0.660
Proposed LRT 0.0061 22.97 24.67 / 0.821
4.4 Ablations
We conducted ablation studies to analyze the impact of key components of RestoreVAR.
Continuous vs. Discrete Conditioning. The RestoreVAR transformer conditions on the continuous latent of the degraded image ($f_{\text{cont}}^{\text{deg}}$). While conditioning with discrete multi-scale latents appears more aligned with VAR’s multi-scale prediction objective, it results in significantly worse performance. To demonstrate this, we train RestoreVAR with discrete and continuous conditioning for $15$ epochs each. As shown in Fig. 8, RestoreVAR with discrete conditioning exhibits much lower validation accuracy.
Discriminator for VAE fine-tuning. As described in Sec. 3.2.2, we fine-tune the VAE decoder on continuous latents using a combination of pixel-level loss and an adversarial loss. To analyze the impact of the discriminator, we compare the reconstructions of VAE decoders fine-tuned with and without the adversarial loss. As shown in Fig. 8, removing the discriminator leads to blurrier reconstruction while including it yields sharper and perceptually better looking outputs.
Latent Refiner Transformer. The Latent Refiner Transformer (LRT) is critical for preserving pixel-level detail in restored images. To analyze its impact, we compare four RestoreVAR variants: (i) No refiner, (ii) HART’s diffusion refiner hart , (iii) LRT without final block outputs, and (iv) our proposed LRT. As shown in Table 5, our LRT achieves the best PSNR and SSIM, while maintaining low inference time and a low parameter count. Using no refiner yields poor PSNR/SSIM scores. Removing the last block outputs significantly reduces performance, indicating its importance as pseudo-continuous guidance for refinement. HART’s MLP diffusion-based refiner performs worse than our proposed LRT while having a significantly higher parameter count and runs $\sim\mathbf{7\times}$ slower.
More ablations are provided in the supplementary.
Figure 7: Validation accuracy of RestoreVAR under discrete vs. continuous conditioning.
Input image
w/o Disc
w Disc
Figure 8: Image reconstructed by VAE decoders fine-tuned on continuous latents with (w) and without (w/o) a discriminator (Disc).
5 Conclusions
We proposed RestoreVAR, a fast and effective generative approach for All-in-One image Restoration (AiOR). Built on the Visual Autoregressive Modeling (VAR) backbone, RestoreVAR benefits from VAR’s strong generative priors and significantly faster inference compared to latent diffusion models (LDMs). To tailor VAR for AiOR, we introduced cross-attention mechanisms that inject semantic information from the degraded image into the generation process. Additionally, we proposed a non-generative latent refiner transformer to convert discrete latents to continuous ones, along with a VAE decoder fine-tuned on continuous latents, which together improve reconstruction fidelity. RestoreVAR achieves state-of-the-art performance among generative AiOR models, significantly outperforming LDM-based methods while delivering over 10×10$\times$10 × faster inference and strong generalization to real-world degradations.
Acknowledgments
This work is supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0076. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.