### v2

# RestoreVAR: Visual Autoregressive Generation for All-in-One Image Restoration

Abstract
The use of latent diffusion models (LDMs) such as Stable Diffusion has significantly improved the perceptual quality of All-in-One image Restoration (AiOR) methods, while also enhancing their generalization capabilities. However, these LDM-based frameworks suffer from slow inference due to their iterative denoising process, rendering them impractical for time-sensitive applications. Visual autoregressive modeling (VAR), a recently introduced approach for image generation, performs scale-space autoregression and achieves comparable performance to that of state-of-the-art diffusion transformers with drastically reduced computational costs. Moreover, our analysis reveals that coarse scales in VAR primarily capture degradations while finer scales encode scene detail, simplifying the restoration process. Motivated by this, we propose RestoreVAR, a novel VAR-based generative approach for AiOR that significantly outperforms LDM-based models in restoration performance while achieving over $\mathbf{10\times}$ faster inference. To optimally exploit the advantages of VAR for AiOR, we propose architectural modifications and improvements, including intricately designed cross-attention mechanisms and a latent-space refinement module, tailored for the AiOR task. Extensive experiments show that RestoreVAR achieves state-of-the-art performance among generative AiOR methods, while also exhibiting strong generalization capabilities.
1 Introduction
Image restoration is a complex inverse problem that aims to recover clean images from degradations, such as haze, rain, snow, blur, and low-light conditions. Recently, the paradigm of All-in-One image Restoration (AiOR) has emerged, where a single network is trained to handle multiple degradation types. Existing AiOR methods can be broadly categorized into non-generative and generative approaches. Non-generative models such as AirNet (Li et al., 2022), PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan & Patel, 2024), and AdaIR (Cui et al., 2024), deterministically map degraded images to their clean counterparts. While these methods offer fast inference and reliable pixel-level restoration performance, they often fail to generalize to diverse degradations encountered in real-world scenarios. To overcome this challenge, recent works have adopted generative models that aim to capture the distribution of clean images and produce more perceptually realistic outputs. Early works (Chen et al., 2022; Kupyn et al., 2018) based on GANs (Goodfellow et al., 2020) attempted this through adversarial learning, but suffered from mode collapse and unstable training. To improve fidelity and training stability, DiffUIR (Zheng et al., 2024) and DA-CLIP (Luo et al., 2023) employed pixel-space diffusion models (Ho et al., 2020). However, their high computational cost makes large-scale pretraining infeasible, limiting their ability to learn strong generative priors. In contrast, recent methods such as Diff-Plugin(Liu et al., 2024), AutoDIR(Jiang et al., 2023), and PixWizard (Lin et al., 2024) leverage latent diffusion models (LDMs), such as Stable Diffusion (Rombach et al., 2022). By operating in a latent space, LDMs significantly reduce computational costs, enabling large-scale pretraining which equips them with strong generative priors of natural images. These priors allow LDM-based AiOR methods to deliver perceptually realistic restoration and improved generalization to real-world degradations.
Despite their advantages, LDM-based AiOR methods have some shortcomings. (1) LDMs require multiple denoising steps during inference, resulting in significantly longer runtimes compared to non-generative models. Their slow inference speeds pose challenges for applications that demand real-time processing, such as video surveillance or autonomous navigation. (2) LDMs rely on variational autoencoders (VAEs) (Kingma, 2013) which are primarily trained for generative diversity, rather than accurate pixel-level reconstruction. Consequently, the restored images obtained from LDM-based AiOR methods exhibit loss of fine structural details, hindering their performance.
Autoregressive models have driven rapid advances in natural language processing through large language models (LLMs) such as GPT-3 (Radford et al., 2019b) and LLaMA (Touvron et al., 2023). These models generate outputs by predicting the next token, conditioned on previously generated tokens. Recently, Visual AutoRegressive (VAR) Modeling (Tian et al., 2024) introduced scale-space autoregression for image generation, performing next-scale prediction in the latent space of a multi-scale vector-quantized VAE (VQVAE). VAR achieves performance comparable to state-of-the-art diffusion models such as DiT-XL/2 (Peebles & Xie, 2023), while operating $45\times$ faster. Despite its success in generative tasks, the application of VAR to low-level vision tasks such as image restoration remains largely unexplored. To the best of our knowledge, only two prior works—VarSR (Qu et al., 2025) and Varformer (Wang & Zhao, 2024)—have used VAR for image restoration. VarSR focused exclusively on the super-resolution task, while Varformer utilized intermediate VAR features to guide a separate non-generative network for AiOR. In contrast, our approach is generative and fully exploits the strong priors of the pretrained VAR model by training it directly for the AiOR task. Our analysis in Sec. 3.2 also reveals that the scale-space decomposition of VAR captures degradations predominantly in coarse scales and scene-level details in fine scales, making it well-suited for AiOR.
(a) Overview
[…]
(c) Model complexity (bubble size $\propto$ parameter count)
Figure 1: RestoreVAR, our proposed VAR-based (Tian et al., 2024) scale-space generative AiOR model (a), significantly outperforms LDM-based methods as shown in (b). RestoreVAR also offers drastic reductions in computational complexity as shown in (c).
To this end, we introduce RestoreVAR, a novel generative approach for AiOR that addresses some of the key limitations of LDM-based approaches. Firstly, RestoreVAR adopts the autoregressive structure of VAR, achieving state-of-the-art generative AiOR performance with over $\mathbf{10\times}$ faster inference than LDM-based methods (see Fig. 1). Secondly, RestoreVAR employs cross-attention mechanisms conditioned on the degraded image latents, enabling the model to maintain spatial consistency and minimize hallucinations. Thirdly, to mitigate the loss of fine details by the vector quantization and VAE decoding processes, we propose a lightweight (only $\sim\mathbf{3}$% overhead) non-generative latent refinement transformer which predicts de-quantized latents from the outputs of VAR. Additionally, we fine-tune the VAE decoder to operate on these continuous latents, further enhancing reconstruction quality. Finally, through extensive experiments, we demonstrate that RestoreVAR achieves state-of-the-art performance among generative restoration models, while also exhibiting strong generalization to real-world degradations. To summarize, our key contributions are:
  1. 1.
    We propose RestoreVAR, the first VAR-based generative AiOR framework that achieves superior performance and a $\mathbf{10\times}$ faster inference than LDM-based methods.
  2. 2.
    To achieve semantically coherent restoration, we introduce degraded image conditioning through cross-attention at each block of the VAR transformer.
  3. 3.
    To mitigate the loss of fine details in the vector quantization and VAE decoding processes, we introduce a non-generative latent refiner transformer which converts discretized latents into continuous ones, and fine-tune the VAE decoder to operate on continuous latents.
  4. 4.
    Extensive experiments show that RestoreVAR attains state-of-the-art performance among generative AiOR approaches, with perceptually preferable results and strong generalization.
2 Related Works
2.1 Image restoration
Early restoration models primarily addressed specific degradations  (He et al., 2009; Zhang et al., 2020; Wang et al., 2019; Yasarla & Patel, 2019; Zhang et al., 2021a; Nah et al., 2017). Other methods such as Restormer (Zamir et al., 2022), MPRNet (Zamir et al., 2021) and SwinIR (Liang et al., 2021) introduced architectures for any single restoration task. However, they are restricted to handle one degradation at a time, making them ineffective for multiple degradations. All-in-One image Restoration (AiOR) methods aim to tackle multiple corruptions with a single model. Early approaches include non-generative models such as All-in-one (Li et al., 2020) and Transweather (Valanarasu et al., 2022). PromptIR (Potlapalli et al., 2024) used learnable prompts while AWRaCLe (Rajagopalan & Patel, 2024) utilized visual in-context learning to extract degradation characteristics. Other approaches such as InstructIR (Conde et al., 2025) adopted textual guidance, and DCPT (Hu et al., 2025) proposed a novel pre-training strategy for AiOR. DFPIR (Tian et al., 2025) proposed a feature perturbation strategy for AiOR. Recent AiOR methods have adopted diffusion models. Pixel-space diffusion models (PSDMs) such as DA-CLIP (Luo et al., 2023) and DiffUIR (Zheng et al., 2024) demonstrated improved AiOR performance but lacked robust generative priors. Recent methods have utilized the strong priors of LDMs for AiOR. Diff-Plugin (Liu et al., 2024) adopts task plugins to guide an LDM for AiOR. AutoDIR (Jiang et al., 2023) automatically detects and restores degradations using an LDM. PixWizard (Lin et al., 2024) is a multi-task SD-XL (Podell et al., 2023) based model capable of performing AiOR among other tasks. However, LDM-based approaches are slow at inference time—a limitation we aim to overcome using visual autoregressive modeling (VAR).
2.2 Autoregressive models in vision
Recent works (Van Den Oord et al., 2016; Tian et al., 2024) have extended autoregressive (AR) models to vision and can be categorized as pixel‐space AR (Van Den Oord et al., 2016; Van den Oord et al., 2016; Chen et al., 2018), token‐based AR (Van Den Oord et al., 2017; Yu et al., 2023; Ramesh et al., 2021) and scale‐space AR (Tian et al., 2024; Ren et al., 2024; Guo et al., 2025). Pixel-space AR predicts raw pixels one by one in raster order, as in PixelRNN (Van Den Oord et al., 2016) and PixelCNN++ (Salimans et al., 2017), but is very slow at high resolutions. Token-based AR compresses images into discrete latent codes via vector quantization (e.g., VQ-VAE (Van Den Oord et al., 2017), VQGAN (Esser et al., 2021)) and then models code sequences with transformers (e.g. ImageGPT (Chen et al., 2020)). This trades-off codebook size and transformer capacity against tractability for high-resolution generation. Scale-space AR, as introduced in VAR (Tian et al., 2024), generates latents from coarse to fine scales and matches the quality of Diffusion Transformers (Peebles & Xie, 2023) at a fraction of the inference cost. HART (Tang et al., 2024) scales VAR to higher resolution and uses a MLP-based diffusion refiner to convert discrete VAR latents into continuous representations. Despite VAR’s success in generative tasks, it remains underexplored for image restoration with only two prior works-VarSR (Qu et al., 2025) and Varformer (Wang & Zhao, 2024). VarSR addressed super-resolution, while Varformer used VAR’s features to guide a non-generative AiOR model. In contrast, RestoreVAR is a generative model which directly trains VAR for AiOR.
3 Proposed Method
We first explain the working principles behind VAR for image generation. We then describe our scale-space analysis of VAR and detail RestoreVAR, our proposed VAR-based approach for AiOR.
3.1 Preliminaries: Visual Autoregressive Modelling
Visual Autoregressive Modelling, or VAR, is a novel autoregressive class-conditioned image generation method which uses a GPT-2 (Radford et al., 2019a) style decoder-only transformer architecture for next-scale prediction. The VAR transformer operates in the latent space of a multi-scale VQVAE which uses $K$ scales. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the VQVAE encoder outputs a latent representation $f_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. Hereafter, we will refer to $f_{\text{cont}}$ as the continuous latent, and the latent obtained after quantization as discrete latent. Instead of directly quantizing $f_{\text{cont}}$, a multi-scale residual quantization using a shared codebook across $K$ spatial scales is performed. First, the residual and accumulated quantized (or discrete) reconstruction of $f_{\text{cont}}$ are initialized as $f_{\text{res}}^{(0)}:=f_{\text{cont}}$ and $f_{\text{quant}}^{(0)}:=0$, respectively. At each scale $k=1,\ldots,K$, an index map $r_{k}\in\mathbb{Z}^{H_{k}\times W_{k}}$ is obtained by quantizing the downsampled residual feature:
$r_{k}:=\text{quantize}\left(\text{downsample}\left(f_{\text{res}}^{(k-1)}\right)\right).\vskip-3.0pt$
The indices $r_{k}$ are then decoded using the codebook embeddings $e(\cdot)$, upsampled to match the full resolution, and refined using a convolutional module $\phi_{k}(\cdot)$ to obtain
$h_{k}:=\phi_{k}\left(\text{upsample}\left(e(r_{k})\right)\right),\in\mathbb{R}^{H_{K}\times W_{K}\times C}.\vskip-3.0pt$
This is done to approximate the information captured at the current scale which is used to update the residual continuous features to be modelled by subsequent scales as
$f_{\text{quant}}^{(k)}:=f_{\text{quant}}^{(k-1)}+h_{k},\quad f_{\text{res}}^{(k)}:=f_{\text{cont}}-f_{\text{quant}}^{(k)}.\vskip-3.0pt$
This process is repeated for all scales and yields a set of index maps $\{r_{1},r_{2},\ldots,r_{K}\}$, each consisting of the code-book indices of residual information at an increasingly finer scale.
For training, VAR uses teacher-forcing, where the ground-truth index maps $\{r_{1},r_{2},\ldots,r_{K}\}$ are used to autoregressively predict the next scale. For each scale $k$, the accumulated reconstruction $f_{\text{quant}}^{(k-1)}=\sum_{i=1}^{k-1}\phi_{i}\left(\text{upsample}(e(r_{i}))\right)$ is interpolated to the resolution of scale $k$ to obtain $\hat{f}_{\text{quant}}^{(k)}$, which is then flattened into tokens, and concatenated with the remaining tokens to form the input sequence. A start-of-sequence (SOS) token, derived from the class label embedding, is then prepended to this input sequence. A block-wise causal attention mask is used to ensure that predictions for scale $k$ attend only to the previous scales. VAR is trained to minimize the cross-entropy loss between predicted logits and the ground-truth index maps, modeling the likelihood
$p(r_{1},r_{2},\ldots,r_{K})=\prod_{k=1}^{K}p(r_{k}\mid r_{1},r_{2},\ldots,r_{k-1}).\vskip-4.0pt\vskip-3.0pt$
During inference, the SOS token is created from the target class label. VAR then autoregressively predicts each index map $r_{k}$, one scale at a time. After predicting $r_{k}$, its embedding is upsampled, refined and accumulated to form the input for the next scale, mimicking the same procedure used during training. The VAR model uses only $K=10$ latent scales with key-value (KV) caching, enabling significantly faster inference compared to latent diffusion models.
3.2 Scale-Space Analysis of VAR
In addition to VAR’s competitive performance to LDMs with far superior inference speed, we found that its residual scale-space decomposition focuses on degradations and scene-level details across different scales. To demonstrate this, we consider clean (GT)–degraded image pairs and compute their scale-wise residual indices $\{r^{\text{GT}}_{k}\}_{k=1}^{K}$ and $\{r^{\text{Deg}}_{k}\}_{k=1}^{K}$, respectively, where $K=10$. We define coarse scales as $k=1,\ldots,5$ (low-resolution index maps) and fine scales as $k=6,\ldots,10$ (higher-resolution index maps). The first and last columns of Fig. 2 show reconstructions from $r^{\text{Deg}}_{k}$ and $r^{\text{GT}}_{k}$, respectively. In column 2, we replace the coarse scales of $r^{\text{GT}}_{k}$ with those from $r^{\text{Deg}}_{k}$. This introduces the degradation, although fine scales remain unchanged. Next, in column 3 we replace the fine scales of $r^{\text{GT}}_{k}$ with those from $r^{\text{Deg}}_{k}$, which yields a clean image with some loss of fine details. These observations indicate that coarse scales in VAR capture degradations, while finer scales encode scene-level detail. Notably, this observation holds across multiple degradations and simplifies restoration for VAR as removing degradations requires correctly predicting only the early scales which contain a small number of tokens, while scene details can be reconstructed in subsequent scales.
3.3 RestoreVAR
Degraded GT++coarse GT++fine GT
Haze
Refer to caption Refer to caption Refer to caption Refer to caption
Snow
Refer to caption Refer to caption Refer to caption Refer to caption
Rain
Refer to caption Refer to caption Refer to caption Refer to caption
Low-light
Refer to caption Refer to caption Refer to caption Refer to caption
Blur
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: VAR captures degradations in early scales (coarse) and scene-level details in later scales (fine). Degraded and GT are VQVAE reconstructions of the degraded and ground truth images. GT++coarse replaces early GT scales with degraded ones, while GT++fine replaces the late GT scales.
We now describe RestoreVAR, our proposed approach that effectively adapts VAR for AiOR, leveraging its substantial inference speed advantage over LDMs. Given a degraded image $I_{\text{deg}}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict a clean output $I_{\text{clean}}$, close to the ground-truth $I_{\text{gt}}$. Adapting VAR to AiOR is non-trivial due to the need for high-quality pixel-level reconstruction, which is compromised by two factors: (1) VAR’s strong generative priors can cause hallucinations in the restored images without proper conditioning. (2) Vector quantization and VAE decoding introduce artifacts that hinder pixel-level restoration. RestoreVAR addresses these challenges through architectural enhancements, including cross-attention to incorporate semantic guidance from the degraded image, and a novel non-generative transformer that refines discrete latents into their continuous form to preserve fine details in the restored image. We describe these components below.
Figure 3: Illustration of RestoreVAR for training and inference. (a) Shows the training procedure for each component of RestoreVAR, and (b) shows the overall pipeline during inference.
3.3.1 Autoregressive Transformer Architecture
For training, the multi-scale teacher-forcing input is constructed from the ground-truth image $I_{\text{gt}}$ (see Sec. 3.1). The start-of-sequence (SOS) token is computed from a fixed label index and augmented with a global context vector derived from the degraded image (see supplementary for details). These features are flattened and concatenated into a token sequence $\hat{f}_{\text{quant}}\in\mathbb{R}^{L\times C}$, where $L$ is the total number of tokens across all scales (see Fig. 3(a)(i)). The VAR transformer is then trained to autoregressively predict the next-scale indices $\{r^{k}_{\text{gt}}\}_{k=1}^{K}\in\mathbb{R}^{L}$ of the clean image.
To enable semantically consistent restoration, we inject information from the degraded image through cross-attention at each transformer block. At block $i$, the queries are given by the output of the feed-forward network ($x_{\text{block}_{i}}\in\mathbb{R}^{L\times D}$, where $D$ is the embedding dimension), while keys and values are derived from the continuous latent of the degraded image, $f_{\text{cont}}^{\text{deg}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. This latent is reshaped into a sequence of conditioning tokens and is appropriately projected to the embedding dimension of the transformer. As shown in Sec. 4.4, conditioning on continuous latents significantly outperforms conditioning on discrete ones. To summarize, cross-attention ($\text{CA}(\cdot,\cdot)$) is applied as
$x_{\text{block}_{\text{CA}}}=x_{\text{block}_{i}}+g_{i}\times\text{CA}(x_{\text{block}_{i}},f_{\text{cont}}^{\text{deg}}).$
We initialize $g_{i}=0$ to retain VAR’s pretrained behavior and gradually introduce conditioning. Furthermore, we replace absolute positional embeddings in VAR with 2D Rotary Positional Embeddings (RoPE) for scaling resolution from $256\times 256$ to $512\times 512$, as RoPE is well-suited for handling varying sequence lengths (Su et al., 2024). We also remove AdaLN layers, reducing $\sim 100$M parameters with negligible impact on performance. Inference closely follows that of VAR (see Sec. 3.1), except that each scale prediction is now guided by the degraded latent. The output is a sequence of predicted indices $\{\hat{r}^{k}_{\text{pred}}\}_{k=1}^{K}$, which is then used to construct the discrete restored latent $f^{\text{pred}}_{\text{quant}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. The above steps are shown in Fig. 3(a)(ii). More architectural details are given in the supplementary.
3.3.2 Detail-Preserving Restoration
The discrete latent ($f^{\text{pred}}_{\text{quant}}$) predicted by the RestoreVAR transformer is decoded by the VQVAE to produce the restored image. However, vector-quantization and VAE decoding cause a noticeable loss of fine details in the pixel-space, leading to distorted reconstructions. This presents a major challenge for using VAR in AiOR, as the scene semantics may not be accurately preserved. To address this, we introduce VAE decoder fine-tuning on continuous latents, and a lightweight latent refinement transformer (LRT) that converts discrete latents to continuous latents for decoding.
VAE Decoder Fine-Tuning. HART (Tang et al., 2024) addressed VAE-induced distortions by fine-tuning the VAE decoder on both discrete and continuous latents. While effective for generative tasks, the VAE decoder of HART produces overly textured outputs, compromising accurate reconstruction (see supplementary). Instead, we fine-tune the decoder only on continuous latents, bypassing the quantizer. The encoder and quantizer are kept frozen, and the decoder is trained on $(f^{\text{gt}}_{\text{cont}},I_{\text{gt}})$ pairs. To avoid overly smooth outputs, we use a PatchGAN (Isola et al., 2017) discriminator (see Sec. 4.4) and optimize the decoder using pixel-wise, perceptual, and adversarial losses as
$\mathcal{L}_{\text{dec}}=\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{percep}}+\lambda_{4}\mathcal{L}_{\text{adv}},\vskip-4.0pt$
where $\mathcal{L}_{\text{L1}}$ is the L1 loss, $\mathcal{L}_{\text{SSIM}}$ is the SSIM loss, $\mathcal{L}_{\text{percep}}$ is the perceptual loss, $\mathcal{L}_{\text{adv}}$ is the adversarial loss and $\lambda_{i}$ are their respective weights (see Fig. 3(a)(iv)). Our fine-tuning approach yields a decoder that is well-aligned with the objectives of AiOR, achieving mean (over $1000$ samples) reconstruction PSNR/SSIM scores of 28.1428.14dB/$0.842$, outperforming both the VAR VQVAE (22.5922.59dB/$0.679$) and HART decoders (26.4826.48dB/$0.804$). Qualitative comparisons are given in the supplementary.
Refining Discrete Latents. Since the VAE decoder is fine-tuned for continuous latents, the predicted discrete latent, $f^{\text{pred}}_{\text{quant}}$, must be converted into a continuous form for decoding.
Input
Discrete
Refiner
Continuous
Figure 4: Illustration of images decoded from discrete and continuous latents, along with the refiner’s predicted residuals.
While HART uses a 3737M parameter diffusion-based MLP for this, it incurs a $\sim 20\%$ inference overhead due to iterative denoising. Instead, we propose a lightweight, non-generative latent refinement transformer (LRT) that predicts a residual, which when added to $f^{\text{pred}}_{\text{quant}}$, produces a continuous latent, $\hat{f}_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$ as
$\hat{f}_{\text{cont}}=f^{\text{pred}}_{\text{quant}}+\text{LRM}(f^{\text{pred}}_{\text{quant}},z),\vskip-4.0pt$
where $z\in\mathbb{R}^{L\times D}$ is the output from the final RestoreVAR transformer block. $z$ is passed through cross-attention and provides pseudo-continuous guidance to the LRT which is critical for performance (see Sec. 4.4). The LRT is trained using $\mathcal{L}_{1}$ loss between the predicted and ground-truth continuous latents ($f_{\text{cont}}^{\text{gt}}$) as $\mathcal{L}_{\text{LRT}}=\mathcal{L}_{1}(\hat{f}_{\text{cont}},f_{\text{cont}}^{\text{gt}})$. Our LRT introduces only $3\%$ additional overhead and significantly outperforms HART’s refiner in PSNR and SSIM scores (see Sec. 4.4). The training procedure of the LRT is shown in Fig. 3(a)(iii) and Fig. 4 provides a visual example of its predictions.
Thus, RestoreVAR combines the VAR transformer, LRT, and fine-tuned decoder to deliver fast, perceptually realistic, and structurally faithful results. Fig. 3(b) depicts inference of RestoreVAR.
4 Experiments
In this section, we provide implementation details, comparisons with existing All-in-One image Restoration (AiOR) approaches, and present ablations on key components of our framework.
4.1 Implementation Details
[…]
4.2 Datasets
We trained RestoreVAR for five tasks: dehazing, desnowing, deraining, low-light enhancement and deblurring. For dehazing, we used the RESIDE (Li et al., 2019) dataset comprising $72135$ training and $500$ test images. The Snow100k dataset (Liu et al., 2018) was used for desnowing, with $50000$ training and $16801$ test images (heavy subset). For deraining, we used Rain13K (Zamir et al., 2021) consisting of $13711$ training and $4298$ test images. The LOLv1 (Wei et al., 2018) dataset was used for low-light enhancement, consisting of $485$ training and $15$ test images. For deblurring, we used the GoPro (Nah et al., 2017) dataset comprising $2103$ training and $1111$ test images. We also assess generalization performance on real-world, unseen and mixed degradation datasets, namely, LHP (Guo et al., 2023) ($1000$ images), REVIDE (Zhang et al., 2021b) ($284$ images), TOLED (Zhou et al., 2021) ($30$ images), POLED (Zhou et al., 2021) ($30$ images), CDD (Guo et al., 2024) ($200$ images, mix of haze and rain), and LOLBlur (Zhou et al., 2022) ($482$ images, mix of low-light and blur). TOLED and POLED datasets contain unseen degradation of under-display camera restoration.
Table 1: Quantitative comparisons of RestoreVAR with the state-of-the-art LDM-based generative AiOR approaches, and non-generative methods. RestoreVAR significantly outperforms generative methods on PSNR, SSIM and LPIPS scores. The best generative approach is indicated in bold.
Method Venue RESIDE Snow100k Rain13K LOLv1 GoPro
[…]
PromptIR NeurIPS’23 32.02 0.952 0.013 31.98 0.924 0.115 29.56 0.888 0.087 22.89 0.847 0.296 27.21 0.817 0.250
InstructIR ECCV’24 26.90 0.952 0.017 29.56 0.885 0.088 22.81 0.836 0.132 28.26 0.870 0.146
AWRaCLe AAAI’25 30.81 0.979 0.013 30.56 0.904 0.088 31.26 0.908 0.068 21.04 0.818 0.146 26.78 0.820 0.248
DCPT ICLR’25 29.10 0.968 0.017 24.11 0.766 0.203 23.67 0.863 0.106 27.92 0.877 0.169
DFPIR CVPR’25 31.39 0.979 0.012 24.87 0.794 0.171 23.12 0.853 0.123 28.66 0.884 0.158
[…]
Input Diff-Plugin AutoDIR PixWizard RestoreVAR GT
RESIDE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rain13K
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GoPro
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparisons of RestoreVAR with LDM-based generative AiOR approaches. RestoreVAR achieves consistent restoration with enhanced preservation of fine-details.
4.3 Comparisons
We compare RestoreVAR with state-of-the-art generative and non-generative methods for AiOR. For non-generative approaches, we include PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan & Patel, 2024), DCPT (Hu et al., 2025) and DFPIR (Tian et al., 2025). Among generative methods, we compare with the LDM-based approaches Diff-Plugin (Liu et al., 2024), AutoDIR (Jiang et al., 2023) and PixWizard (Lin et al., 2024). To ensure a fair comparison, we retrained PromptIR and AWRaCLe, as their official checkpoints were not trained for most of our AiOR tasks. All other methods were evaluated using their publicly released checkpoints. For AutoDIR, we report results without the structure correction module, as this module functions as an independent, non-generative restoration network (more details in supplementary). The results reported for PixWizard were obtained using its publicly released checkpoint. We do not compare with task-specific restoration models, as RestoreVAR is proposed for the AiOR setting.
Table 4.2 presents PSNR, SSIM and LPIPS scores on the RESIDE, Snow100k, Rain13K, LOLv1 and GoPro datasets. RestoreVAR surpasses LDM-based AiOR methods at a fraction of their computational cost (inference time (s) per image)—Diff-Plugin:$2.04$s, AutoDIR: $8.477$s, PixWizard: $8.247$s and RestoreVAR: $0.201$s, highlighting the efficacy of our framework. More detailed complexity comparisons are given in the supplementary along with a derivation showing that the time complexity of VAR with maximum latent resolution $n\times n$ is $\mathcal{O}(\log n)$ lower than an LDM operating at the same latent resolution. Qualitative comparisons with LDM-based methods in Fig. 5 further illustrate that RestoreVAR produces restored images of high quality while better preserving fine details. Visual results for the Snow100k and LOLv1 datasets are provided in the supplementary. While non-generative methods achieve better scores, it is important to recognize that the performance of RestoreVAR is inherently influenced by the quality of the VAE decoder; a limitation shared by all latent generative approaches. Despite this constraint, RestoreVAR narrows the gap with non-generative methods while maintaining the benefits of a generative framework, i.e., perceptually realistic results and strong generalization capabilities. To demonstrate these strengths, we evaluate generalization using no-reference image quality metrics (following prior works (Liu et al., 2024; Jiang et al., 2023; Rajagopalan & Patel, 2024)), and assess perceptual realism through a user study.
Table 2: Quantitative comparisons of RestoreVAR against state-of-the-art non-generative approaches on real-world, unseen and mixed degradations. The best result is indicated in bold.
Method LHP REVIDE TOLED POLED LOLBlur (L ++ B) CDD (H ++ R) Average
[…]
PromptIR 56.780 0.366 61.191 0.459 43.218 0.281 34.536 0.303 33.693 0.166 65.895 0.483 49.219 0.343
InstructIR 58.269 0.359 63.116 0.416 44.985 0.298 23.317 0.241 40.221 0.202 65.491 0.482 49.900 0.333
AWRaCLe 57.889 0.333 59.287 0.368 44.670 0.285 40.533 0.332 38.186 0.171 66.253 0.484 51.470 0.329
DCPT 58.044 0.372 60.011 0.446 44.062 0.314 38.138 0.345 37.393 0.175 68.440 0.544 51.681 0.366
DFPIR 56.483 0.330 61.009 0.450 43.820 0.276 35.668 0.289 36.277 0.163 54.408 0.349 47.611 0.310
[…]
Input InstructIR AWRaCLe DCPT DFPIR RestoreVAR GT
LHP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
REVIDE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
TOLED
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
POLED
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CDD
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative comparisons of RestoreVAR with non-generative methods on real, unseen and mixed degradations. RestoreVAR consistently achieves better results.
For testing generalization, we report MUSIQ (Ke et al., 2021) and CLIPIQA (Wang et al., 2023) scores in Table 2 on the real-world, unseen and mixed degradation datasets discussed in Sec. 4.2.
Table 3: Mean scores from user study.
[…]
PromptIR 2.11
InstructIR 2.93
AWRaCLe 2.33
DCPT 2.42
DFPIR 2.35
AutoDIR 3.68
[…]
RestoreVAR achieves higher scores than non-generative models (on average), indicating better robustness under these degradations. Qualitative results for this experiment are shown in Fig. 6, where RestoreVAR consistently outperforms non-generative approaches. Due to space constraints, qualitative comparisons with PromptIR and visual results for LOLBlur are given in the supplementary. To further evaluate perceptual quality, we conducted a user study in which participants rated outputs from non-generative models, AutoDIR (LDM-based) and RestoreVAR, for $50$ real-world scenes. We received $36$ responses with each participant scoring outputs based on scene consistency, restoration quality, and overall appeal on a 5-point scale. Table 3 shows that RestoreVAR received the highest average ratings (across all three criteria), highlighting its ability to produce images that align closely with human preferences.
These results highlight the effectiveness of RestoreVAR for AiOR. We discuss limitations of RestoreVAR in the supplementary.
4.4 Ablations
Continuous vs. Discrete Conditioning. The RestoreVAR transformer conditions on the continuous latent of the degraded image ($f_{\text{cont}}^{\text{deg}}$). While conditioning with discrete multi-scale latents appears more aligned with VAR’s multi-scale prediction objective, it results in significantly worse performance. To demonstrate this, we train RestoreVAR with discrete and continuous conditioning for $15$ epochs each. As shown in Fig. 8, RestoreVAR with discrete conditioning exhibits much lower validation accuracy.
Discriminator for VAE fine-tuning. As described in Sec. 3.3.2, we fine-tune the VAE decoder on continuous latents using a combination of pixel-level loss and an adversarial loss. To analyze the impact of the discriminator, we compare the reconstructions of VAE decoders fine-tuned with and without the adversarial loss. As shown in Fig. 8, removing the discriminator leads to blurrier reconstruction while including it yields sharper and perceptually better looking outputs.
Table 4: Ablations on the types of latent refiners. Our proposed latent refiner transformer (LRT) performs best, with minimal overhead.
Refiner Type Time (s) Params (M) PSNR / SSIM
No Refiner 21.71 / 0.690
HART Refiner 0.0455 36.06 23.48 / 0.777
LRT w/o Last-Block 0.0036 14.61 21.23 / 0.660
Proposed LRT 0.0061 22.97 24.67 / 0.821
Latent Refiner Transformer. The Latent Refiner Transformer (LRT) is critical for preserving pixel-level detail in restored images. To analyze its impact, we compare four RestoreVAR variants: (i) No refiner, (ii) HART’s diffusion refiner, (iii) LRT without final block outputs, and (iv) our proposed LRT. As shown in Table 4, our LRT achieves the best PSNR and SSIM, while maintaining low inference time and a low parameter count. Using no refiner yields poor PSNR/SSIM scores. Removing the last block outputs significantly reduces performance, indicating its importance as pseudo-continuous guidance for refinement. HART’s MLP diffusion-based refiner performs worse than our LRT while having a much higher parameter count and runs $\sim\mathbf{7\times}$ slower.
More ablations are provided in the supplementary.
Figure 7: Validation accuracy of RestoreVAR under discrete vs. continuous conditioning.
Input image
w/o Disc
w Disc
Figure 8: Image reconstructed by VAE decoders fine-tuned on continuous latents with (w) and without (w/o) a discriminator (Disc).
5 Conclusions
We proposed RestoreVAR, a fast and effective generative approach for AiOR. Built on the VAR backbone, RestoreVAR benefits from VAR’s strong generative priors and significantly faster inference compared to LDMs. To tailor VAR for AiOR, we introduced cross-attention mechanisms that inject semantic information from the degraded image into the generation process. Additionally, we proposed a non-generative latent refiner transformer to convert discrete latents to continuous ones, along with a VAE decoder fine-tuned on continuous latents, which together improve reconstruction fidelity. RestoreVAR achieves state-of-the-art performance among generative AiOR models, outperforming LDM-based methods while delivering over $10\times$ faster inference and strong generalization.
Acknowledgments
This work is supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0076. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.